# Impact of AI Search Summaries on Website Traffic: Evidence from Google AI Overviews and Wikipedia

### v1

@@ 2 Setting and Data
We then restrict attention to comparable cross-language units by retaining only articles that (i) have an English version and (ii) have at least one corresponding version in the control languages (Hindi, Japanese, Portuguese, or Indonesian). In the resulting analysis sample, we observe English article pages and total article-language pairs across all five languages (Table 3). For each retained article-language pair, we query the Wikimedia Analytics API for daily page views over our analysis window surrounding the March 22, 2024 AIO test start, and assemble these into an article-language-day panel.
@@ 4 Economic Impact and Implications
That is equivalent to an annual loss of income of M USD for a publisher like Wikipedia. To the extent that the traffic and ad revenues lost by publishers like Wikipedia are transferred to AIOs (and ads shown within the AIOs), these numbers can be interpreted as a transfer in ad revenue from publishers to search engines.
@@ 5 Conclusion
We find clear evidence of traffic displacement. After the introduction of Google’s AIO feature in the US, English Wikipedia page views fell by about 15% relative to the comparison languages. Effects are heterogeneous across topics: declines are largest for Culture and smallest for STEM, consistent with stronger substitution when a short synthesized answer can resolve the query. Because exposure is imperfect in a language-based design (global English readership and gradual rollout), we interpret these estimates as conservative lower bounds. Our findings imply that generative-answer features such as AIO in search engines can materially reallocate attention away from publisher pages, with direct consequences for monetization via advertising and the top of the funnel for subscriptions and commerce.
We close with a few caveats and suggestions for future research. First, our research focuses on one large publisher (Wikipedia); it is possible that the magnitude and direction of the effect of AIO on other publishers is different. In particular, we note that Wikipedia is unique in two ways. On the one hand, it is the canonical answer repository for informational queries, for which AIOs serve as a direct substitute. On the other hand, it is also one of the websites that tends to be heavily cited in AIOs (unlike many smaller publishers), which may have mitigated the impact on Wikipedia. Thus, it is hard to extrapolate the effect to other types of websites. It would be interesting to see how AIOs and generative summaries in search engines affect traffic to smaller publishers and/or sites that are largely transactional or e-commerce-based. Second, while the focus here was on Google AIO, our findings are also relevant to other traditional search engines such as Bing, as well as AI-based answer engines or summarization tools; for example, OpenAI’s ChatGPT Search (25) and Perplexity AI (26). Thus, the overall effect on publishers can be amplified even further. Finally, AIOs and LLM-based summarization features are still evolving, and their long-term impact on the digital economy is still unclear. One interesting direction would be to see how publisher quality and incentives to contribute to repositories such as Wikipedia changes, and how this, in turn, affects AIOs themselves.

### v3

@@ 2 Setting and Data
We then restrict attention to comparable cross-language units by retaining only articles that (i) have an English version and (ii) have at least one corresponding version in the control languages (Hindi, Japanese, Portuguese, or Indonesian). In the resulting analysis sample, we observe English article pages and total article-language pairs across all five languages (Table 3). For each retained article-language pair, we query the Wikimedia Analytics API for daily page views over our analysis window surrounding the March 22, 2024, AIO test start, and assemble these into an article-language-day panel.
@@ 4 Economic Impact and Implications
That is equivalent to an annual loss of income of M USD for a publisher like Wikipedia. To the extent that the traffic and ad revenues lost by publishers like Wikipedia are transferred to AIOs (and ads shown within the AIOs), these numbers can interpreted as a transfer in ad revenue from publishers to search engines.
@@ 5 Conclusion
We find clear evidence of traffic displacement. After the introduction of Google’s AIO feature in US, English Wikipedia page views fall by about 15% relative to the comparison languages. Effects are heterogeneous across topics: declines are largest for Culture and smallest for STEM, consistent with stronger substitution when a short synthesized answer can resolve the query. Because exposure is imperfect in a language-based design (global English readership and gradual rollout), we interpret these estimates as conservative lower bounds. Our findings imply that generative-answer features such as AIO in search engines can materially reallocate attention away from publisher pages, with direct consequences for monetization via advertising and the top of the funnel for subscriptions and commerce.
We close with a few caveats and suggestions for future research. First, our research focuses on one large publisher (Wikipedia); it is possible that the magnitude and direction of the effect of AIO on other publishers is different. In particular, we note that Wikipedia is unique in two ways. On the one hand, it is the canonical answer repository for informational queries, for which AIOs serve as a direct substitute. On the other hand, it is also one of the websites that tends to be heavily cited in AIOs (unlike many smaller publishers), which may have mitigated the impact on Wikipedia. Thus, it is hard to extrapolate the effect to other types of websites. It would be interesting to see how AIOs and generative summaries in search engines affect traffic to smaller publishers and/or sites that are largely transactional or e-commerce based. Second, while the focus here was on Google AIO, our findings are also relevant to other traditional search engines such as Bing, as well as AI-based answer engines or summarization tools; for example, OpenAI’s ChatGPT Search (25) and Perplexity AI (26). Thus, the overall effect on publishers can be amplified even further. Finally, AIOs and LLM-based summarization features are still evolving and their long-term impact on the digital economy is still unclear. One interesting direction would be to see how publisher quality and incentives to contribute to repositories such as Wikipedia changes, and how this in turn affects AIOs themselves.